How many moles of CaCl2 will be produced from 3.572 g of Ca(OH)2 reacting with HCl?
1
Expert's answer
2020-04-20T15:07:36-0400

Ca(OH)2 + 2HCl = CaCl2 + H2O 3.572/74= 0.048 moles